OpenStudy (anonymous):

I need help rationalizing the numerator of: (1 + cx^2-sqrt(1+x^2))/x^4 The intent is to evaluate the limit as x approaches 0, & to find all values of c for which the limit exists

OpenStudy (anonymous):

\[\frac{1 + cx^2-\sqrt{1+x^2}}{x^4}\] like that?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

\[\frac{(1 + cx^2)-\sqrt{1+x^2}}{x^4}\times \frac{(1 + cx^2)+\sqrt{1+x^2}}{(1 + cx^2)+\sqrt{1+x^2}}\]

OpenStudy (anonymous):

\[\frac{(1+cx^2)^2-(1+x^2)}{x^4\left (1 + cx^2)+\sqrt{1+x^2}\right)}\]

OpenStudy (anonymous):

multiply out, leave the denominator factored, get \[\frac{1+2cx^2+c^2x^4-1-x^4}{x^4\left ((1 + cx^2)+\sqrt{1+x^2}\right)}\] \[=\frac{2cx^2+(c^2-1)x^4}{x^4\left ((1 + cx^2)+\sqrt{1+x^2}\right)}\]

OpenStudy (anonymous):

i guess you can factor a \(x^2\) from the numerator and cancel, not sure what else you can do

OpenStudy (anonymous):

there's no way to get rid of \[\sqrt{1+x^2}\]

OpenStudy (anonymous):

without undoing what we've just done? rationalizing the numerator was the Professor's hint on how to evaluate the limit, but this looks counterproductive

OpenStudy (anonymous):

you can get rid of it in the numerator, but it shows up in the denominator

OpenStudy (anonymous):

that's what I thought, that's why is seems pointless. Thanks for helping me out, I had the algebra a little messed up

OpenStudy (anonymous):

i tried playing with a couple different values of \(c\) you can try as well 1 does not work, .1 does not, but .5 does http://www.wolframalpha.com/input/?i=%281+%2B.5x^2-sqrt%281%2Bx^2%29%29%2Fx^4 lets see if there is more algebra to be done

OpenStudy (anonymous):

it appears c must be <1

OpenStudy (anonymous):

ok a little help from worlfram shows me that the only \(c\) that works is \(\frac{1}{2}\)

OpenStudy (anonymous):

I see that.

OpenStudy (anonymous):

so maybe i messed up with the algebra, because i think the numerator should therefore be \(2c-1)+\text{blah}x^2\)

OpenStudy (anonymous):

I think the algebra is correct.

OpenStudy (anonymous):

if you use .01 for x, and .5 for c, it shows the limit approaching .12

OpenStudy (anonymous):

ok in the numerator we get \[(2c-1)x^2+c^2x^4\] so at least there is the \(2c-1\) term we need

OpenStudy (anonymous):

wait, which step did you correct?

OpenStudy (anonymous):

yes, it is \(\frac{1}{8}\) http://www.wolframalpha.com/input/?i=limit+x+to+0+%281+%2B.5x^2-sqrt%281%2Bx^2%29%29%2Fx^4 look at the series expansion on the bottom here you will see why it only works if \(c=\frac{1}{2}\) http://www.wolframalpha.com/input/?i=limit+x+to+0+%281+%2Bcx^2-sqrt%281%2Bx^2%29%29%2Fx^4

OpenStudy (anonymous):

ok here is my mistake \[\frac{1+2cx^2+c^2x^4-1-x^4}{x^4\left ((1 + cx^2)+\sqrt{1+x^2}\right)}\] should be \[\frac{1+2cx^2+c^2x^4-1-x^2}{x^4\left ((1 + cx^2)+\sqrt{1+x^2}\right)}\]

OpenStudy (anonymous):

numerator is \[\frac{(2c-1)x^2+c^2x^4}{x^4\left ((1 + cx^2)+\sqrt{1+x^2}\right)}\]

OpenStudy (anonymous):

now i can see it, i made a silly mistake before

OpenStudy (anonymous):

\(2c-1\) must be zero, otherwise you have an \(x^2\) term in the numerator which will mess everything up sorry about the algebra goof, i was bush league mistake

OpenStudy (anonymous):

it's ok, I was struggling with this one really bad, I really appreciate your help

OpenStudy (anonymous):

no problem